The equation of the straight line is parallel to the straight line 2y=3x-1


1
Expert's answer
2021-04-14T07:08:13-0400

Solution.


"2y=3x-1, \newline
y=\frac{3}{2}x-\frac{1}{2}."

The equation of the straight line is parallel to the straight line 2y=3x-1 is:


"y=\frac{3}{2}x+c,"

where "c" is some real number. For a example,

"y=\frac{3}{2}x" is parallel to"2y=3x-1."